• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Ole Hansen

    Log entry time 16:10:19 on April25,2004

    Entry number 119278

    keyword=Shift summary

    Shift started with beam dump diffuser water flow problem. This was fixed at 10:30am. The rest of the shift was basically very productive, except for 30minutes lost around lunchtime when MCC took the beam away for tests of the injector. Hall C is unhappy with the beam quality and asked us if we could give up our energy lock, but we refused since the energy spread is crucial for the experiment.